OpenStudy (anonymous):

the displacement of a particle from a point O is s meters at time t seconds after some initial time where s =2t^3 -7t^2+4t+5 what is the greatest velocity in the negative direction?

OpenStudy (anonymous):

s =2t^3 -7t^2+4t+5 s'=6t^2-14t+4 :velocity s''=12t-14 0=12t-14 t=14/12 :time for greatest velocity s'=6(14/12)^2-14(14/12)+4 s'= -4.167

OpenStudy (anonymous):

how come in the solution book it says the answer is 4.25cm/s?

OpenStudy (anonymous):

why is it 4.25.... @.@ but i think i should be 4.167 since it is asking for -s'